OpenStudy (anonymous):

3x+2y=18 find the y intercept

OpenStudy (zpupster):

subtract both sides by 3x then divide both sides by 2 y=mx+b b is the y-intercept

OpenStudy (zpupster):

you aregoing from standard form Ax+By=C to slope-intercept form y=mx+b

OpenStudy (zpupster):

go ahead i will help

OpenStudy (anonymous):

So I got y=18-3x all over 2

OpenStudy (zpupster):

that is right but it is better to divide both terms by 2 go ahead

OpenStudy (zpupster):

can you get it???

OpenStudy (anonymous):

Yea it's 9 and -3x/2

OpenStudy (zpupster):

yes 9 is our intercept when x is 0 y =9 good job!!

OpenStudy (anonymous):

Ohhhh okay I get it thanks